\begin{thebibliography}{12}

\bibitem{Wops}
A.Morozov and Sh.Shakirov,
%\emph{Generation of Matrix Models by $\hat W$-operators},
JHEP, {\bf 0904} (2009) 064, arXiv:0902.2627

\bibitem{BEMS}
G.Borot, B.Eynard, M.Mulase and B.Safnuk,
\emph{A Matrix Model for Simple Hurwitz Numbers and
Topological Recursion},
arXiv:0906.1206v1

\bibitem{KniUFN}
V.Knizhnik,
%{\it Multiloop Amplitudes in the Theory of Quantum Strings
%and Complex Geometry},
Usp. Phys. Nauk, {\bf 159} (1989) 401-453

\bibitem{UFN2} A.Morozov, {\it String Theory, What is it?},
Sov. Phys. Usp. \textbf{35} (1992) 671-714

\bibitem{intau}
I.Krichever, Russ. Math. Surveys {\bf 32} (1977) v.6 p.185
(Usp.Mat.Nauk {\bf 32} (1977) v.6 p.183); \\
M.Jimbo, T.Miwa and M.Sato, Publ. RIMS
{\bf 14} (1978) 223, {\bf 15} (1979) 201, 577, 871, 1531; \\
R.Hirota, Topics in Current Physics {\bf 17} (1980) 157; \\
B.Dubrovin, Sov.Math.Usp. {\bf 36} (1981) v.2 p.11-80; \\
T.Shiota, {\it Characterization of Jacobian Varieties in Terms
of Soliton Equations}, Harvard University, 1984; \\
M. Jimbo, T. Miwa,
%{\it Solitons and Infinite dimensional Lie
%algebras},
Publ.RIMS, Kyoto Univ,. {\bf 19} (1983) 943-1001; \\
K.Ueno, K.Takasaki,
%{\it Toda lattice hierarchy},
Adv.Studies in Pure Math., {\bf 4} (1984) 1-95; \\
M.Mulase, J.Diff.Geom. {\bf 19} (1984) 403-430;
Invent.Math. {\bf 92} (1988) 1-46; \\
A. Mironov, A. Morozov and L. Vinet, Theor.Math.Phys. {\bf 100}
(1995) 890-899,
%(Teor.Mat.Fiz. {\bf 100} (1994) 119-131),
hep-th/9312213; \\
A.Gerasimov, S.Khoroshkin, D.Lebedev, A.Mironov and A.Morozov,
%\emph{Generalized Hirota Equations and Representation Theory.
%I. The case of $SL(2)$ and $SL_q(2)$},
Int.J.Mod.Phys. \textbf{A10} (1995) 2589-2614, hep-th/9405011;\\
I.K.Kostov, I.Krichever, M.Mineev-Weinstein, P.Wiegmann and A.Zabrodin,
%\emph{$\tau$-function for analytic curves},
hep-th/0005259

\bibitem{UFN3}
A.Morozov,
%{\it Integrability and Matrix Models},
Phys.Usp. \textbf{37 }(1994) 1-55, hep-th/9303139;
%{\it Matrix Models as Integrable Systems},
hep-th/9502091;
%\emph{Challenges of matrix models},
hep-th/0502010; \\
A.Mironov,
%{\it 2d gravity and matrix models. I. 2d gravity},
Int.J.Mod.Phys. {\bf A9} (1994) 4355, hep-th/9312212;
%{\it Matrix models of two-dimensional gravity},
Phys.Part.Nucl. {\bf 33} (2002) 537

\bibitem{SW} N.Seiberg and E.Witten,
Nucl.Phys. {\bf B426} (1994) 19-52, hep-th/9407087;
{\bf B431} (1994) 484-550, hep-th/9408099;\\
A.Gorsky, I.Krichever, A.Marshakov, A.Mironov, A.Morozov,
Phys.Lett. {\bf B355} (1995) 466, hep-th/9505035; \\
R.Donagi and E.Witten,
%\emph{Supersymmetric Yang-Mills Systems And Integrable Systems},
Nucl.Phys. {\bf B460} (1996) 299-334, hep-th/9510101; \\
H.Itoyama and A.Morozov,
%\emph{Integrability and Seiberg-Witten Theory: Curves and Periods},
Nucl.Phys. {\bf B477} (1996) 855-877, hep-th/9511126;
%\emph{Prepotential and the Seiberg-Witten Theory},
{\bf B491} (1997) 529-573, hep-th/9512161; hep-th/9601168; \\
A.Gorsky, A.Marshakov, A.Mironov and A.Morozov,
%\emph{N=2 Supersymmetric QCD and Integrable Spin Chains:
%Rational Case $N_f < 2N_c$},
Phys.Lett. B380 (1996) 75-80, hep-th/9603140;
%\emph{A Note on Spectral Curve for the Periodic Homogeneous
%$XYZ$-Spin Chain},
hep-th/9604078;
%\emph{RG Equations from Whitham Hierarchy,}
Nucl.Phys. B527 (1998) 690-716, hep-th/9802007; \\
A.Marshakov, A.Mironov and A.Morozov,
%\emph{ WDVV-like equations in N=2 SUSY Yang-Mills Theory},
Phys.Lett. B389 (1996) 43-52, hep-th/9607109;
%\emph{WDVV Equations from Algebra of Forms},
Mod.Phys.Lett. A12 (1997) 773-788, hep-th/9701014;
%\emph{More Evidence for the WDVV Equations in
%N=2 SUSY Yang-Mills Theories},
Int.J.Mod.Phys. A15 (2000) 1157-1206, hep-th/9701123; \\
N.Nekrasov,
%\emph{Five Dimensional Gauge Theories and Relativistic Integrable Systems},
Nucl.Phys. B531 (1998) 323-344, hep-th/9609219; \\
A.Marshakov, M.Martellini and A.Morozov,
%\emph{Insights and puzzles from branes: 4-D SUSY Yang-Mills from 6-D models},
Phys.Lett. \textbf{B418} (1998) 294-302, hep-th/9706050; \\
H.Braden, A.Marshakov, A.Mironov and A.Morozov,
%\emph{Seiberg-Witten theory for a non-trivial compactification
%from five to four dimensions},
Phys.Lett. B448 (1999) 195-202, hep-th/9812078;
%\emph{The Ruijsenaars-Schneider Model in the Context
%of Seiberg-Witten Theory},
Nucl.Phys. B558 (1999) 371-390, hep-th/9902205;
%\emph{WDVV equations for 6d Seiberg-Witten theory and bi-elliptic curves},
hep-th/0606035

\bibitem{DV}
R.Dijkgraaf, C.Vafa,
%\emph{Matrix Models, Topological Strings, and Supersymmetric Gauge Theories},
Nucl.Phys. \textbf{B644} (2002) 3-20, hep-th/0206255;
%\emph{On Geometry and Matrix Models},
%Nucl.Phys. \textbf{B644} (2002)
21-39, hep-th/0207106;
%\emph{A Perturbative Window into Non-Perturbative Physics},
hep-th/0208048; \\
L.Chekhov and A.Mironov,
%\emph{Matrix models vs. Seiberg-Witten/Whitham theories},
Phys.Lett. \textbf{B552} (2003) 293-302, hep-th/0209085; \\
H.Itoyama and A.Morozov,
%\emph{The Dijkgraaf-Vafa prepotential
%in the context of general Seiberg-Witten theory},
Nucl.Phys.\textbf{B657} (2003) 53-78, hep-th/0211245;
%\emph{Experiments with the WDVV equations for the gluino-condensate
%prepotential: the cubic (two-cut) case},
Phys.Lett. \textbf{B555} (2003) 287-295,
hep-th/0211259;
%\emph{Calculating Gluino-Condensate Prepotential},
Prog.Theor.Phys. \textbf{109} (2003) 433-463, hep-th/0212032;
%\emph{Gluino-Condensate (CIV-DV) Prepotential
%from its Whitham-Time Derivatives},
Int.J.Mod.Phys. \textbf{A18} (2003) 5889-5906,
hep-th/0301136

\bibitem{Hur}
A.Hurwitz,
%{\it Uber Riemann'sche Flachen mit gegebenen Verzweigungpunkten},
Math.Ann. {\bf 39} (1891) 1-61;
%{\it Uber die Anzal der Riemann'sche Flachen mit
%gegebenen Verzweigungpunkten},
Math.Ann. {bf 55} (1902) 51-60; \\
R.Dijkgraaf,
%{\it Mirror symmetry and elliptic curves},
In: {\sl The moduli spaces of curves},
Progress in Math., 129 (1995), 149-163, Brikh\"auser; \\
R.Vakil, {\sl Enumerative geometry of curves via degeneration methods},
Harvard Ph.D. thesis (1997); \\
I.Kostov, M.Staudacher and T.Wynter,
%\emph{Complex Matrix Models and Statistics of Branched Coverings of 2D Surfaces},
Commun.Math.Phys. 191 (1998) 283-298, hep-th/9703189; \\
I.Goulden and D.Jackson,
%\emph{Transitive factorisations into transpositions and holomorphic
%mappings on the sphere},
Proc.Amer.Math.Soc. \textbf{125} (1997) 51-60,
math/9903094; \\
S.Lando and D.Zvonkine,
%{\it On multiplicitites of the
%Lyashko-Looijenga mapping on the discriminant strata},
Funk.Anal.Appl. {\bf 33} 3 (1999) 178-188;
%{\it Counting ramified coverings and intersection theory
%on spaces of rational functions. I},
math.AG/0303218; \\
S.Natanzon and V.Turaev,
%{\it A compactification of
%Hurwitz space},
Topology, {\bf 38} (1999) 889-914; \\
D.Goulden, D.M.Jackson and A.Vainshtein,
%{\it The number of ramifified coverings of the sphere by torus and
%surfaces of higher genera},
Ann. of Comb. 4(2000), 27-46, Brikh\"auser; \\
A.Okounkov,
%{\it Toda equations for Hurwitz numbers},
Math.Rev.Lett. {\bf 7} (2000) 447-453, math/0004128; \\
A.Givental,
%{\it Gromov-Witten invariants and quantization of
%quadratic hamiltonians},
math/0108100; \\
T.Ekedahl, S.Lando, M.Shapiro, A.Vainshtein,
%{\it Hurvitz naumbers and intersections on moduli spaces of curves},
Invent.Math.146(2001),297-327; \\
S.Lando,
%{\it Ramified coverings of the two-dimensional sphere
%and intersection theory in spaces of meromorphic functions
%on algebraic curves},
Russ.Math.Surv., {\bf 57} (2002) 463-533; \\
A.Alexeevski and S.Natanzon,
%{\it Noncommutative two-dimensional
%topological field theories and Hurwitz numbers for real algebraic
%curves},
Selecta Math., New ser. {\bf 12}:3 (2006) 307-377, math.GT/0202164;
%{\it Hurwitz numbers for regular coverings of surfaces
%by seamed surfaces and Cardy-Frobenius algebra of finite groups},
Amer.Math. Soc.Transl. {\bf 224} (2) (2008) 1-25;
Izvestia RAN, {\bf 12}:4 (2008) 3-24; \\
S.Natanzon, Russian Math.Survey {\bf 61}:4 (2006) 185-186;
%{\it Disk single Hurwitz numbers},
arXiv:0804.0242; \\
J.Zhou,
{\it Hodge integrals, Hurwitz numbers and symmetric groups},
math.AG/0308024; \\
A.Okounkov and R.Pandharipande,
%{\it Gromov-Witten theory,
%Hurwitz theory, and completed cycles},
Ann. of Math. {\bf 163} (2006) 517,
math.AG/0204305; \\
T.Graber and R.Vakil,
%{\it Hodge integrals and Hirwitz numbers
%via virtual localization},
Compositio Math.,
{\bf 135} (2003) 25-36; \\
M.Kazarian and S.Lando,
%{\it Towards the intersection theory
%of Hurwitz spaces},
math.AG/0410388;
%{\it An algebro-geometric proof of Witten's conjecture},
math/0601760; \\
M.Kazarian,
%\emph{KP hierarchy for Hodge integrals}, based on the talk at the
%Moscow Workshop on Combinatorics of moduli spaces, Hurwitz numbers and
%cluster algebras (June 2008),
arXiv:0809.3263; \\
S.Lando,
{\it Combinatorial Facets of Hurwitz numbers},
{\sl Appl.of Group Theory to Combinatorics},
%In: {\sl Applications of Group Theory to Combinatorics},
%Koolen, Kwak and Xu, Eds. Taylor \& Francis Group,
%London,
2008, 109-132; \\
V.Bouchard and M.Marino,
\emph{Hurwitz numbers, matrix models and enumerative geometry},
%In: {\sl From Hodge Theory to Integrability and tQFT: tt*-geometry},
%Proceedings of Symposia in Pure Mathematics, AMS (2008),
arXiv:0709.1458;\\
A.Mironov and A.Morozov,
%\emph{Virasoro constraints for Kontsevich-Hurwitz partition function},
JHEP \textbf{0902} (2009) 024, arXiv:0807.2843; \\
A.Mironov, A.Morozov and S.Natanzon,
%\emph{Complete Set of Cut-and-Join Operators in Hurwitz-Kontsevich Theory},
arXiv:0904.4227;
\emph{Integrability and ${\cal N}$-point Hurwitz Numbers}%,
%to appear


\bibitem{fref}
L.Alvarez-Gaume, J.Bost, G.Moore, P.Nelson and C.Vafa,
%{\it Bosonisation on Higher Genus Riemann Surfaces},
Comm.Math.Phys. {\bf 112} (1987) 503;
%{\it Modular forms and the cosmological constant},
Phys.Lett. {\bf 178B} (1986) 41; \\
E.Verlinde and H.Verlinde,
%{\it Chiral Bosonization, Determinants and the String Partition Function},
Nucl. Phys. {\bf B288} (1987) 357-396; \\
E.D'Hoker and D.Phong, Nucl.Phys. {\bf B292} (1987) 109;
%{\it The Geometry of String Perturbation Theory},
Rev.Mod.Phys., {\bf 60} (1988) 917-1065;\\
D.Lebedev and A.Morozov,
%{\it Statistical sums of strings on hyperelliptic surfaces},
Nucl.Phys. {\bf B302} (1988) 163; \\
A.Morozov,
%{\it Two-loop statsum of superstring},
Nucl.Phys. {\bf B303} (1988) 343;\\
%Sov.J.Nucl.Phys. {\bf 48} (1988) 555-564
A.Gerasimov, A.Marshakov, A.Morozov,
M.Olshanetsky and S.Shatashvili,
%{\it Wess-Zumino-Witten model as a theory of free fields},
Int.J.Mod.Phys. {\bf A5} (1990) 2495; \\
%-2589, sec.5; \\
A.Morozov and A.Perelomov,
{\it Strings and Complex Geometry},
%in Modern Problems of Mathematics, VINITI, Moscow, 1990;
Enc.of Math.Sc., {\bf 54} (1993) 197-280; \nn\\
H.Braden, A.Mironov and A.Morozov,
%\emph{QCD, Wick's Theorem for KdV $\tau$-functions and the String Equation},
Phys.Lett. {\bf B514} (2001) 293-298, hep-th/0105169

\bibitem{Mth}
A.Alexandrov, A.Mironov and A.Morozov,
%{\it M-theory of matrix models},
Theor.Math.Phys. {\bf 150} (2007) 179-192, hep-th/0605171;
%{\it Instantons and merons in matrix models},
Physica {\bf D 235} (2007) 126-167, hep-th/0608228;
%{BGW Model  as the Second Constituent of Complex Matrix Model},
arXiv:0906.3305  \\
N.Orantin,
%{\it Symplectic invariants, Virasoro constraints and Givental decomposition },
arXiv:0808.0635


\bibitem{GKM}
M. Kontsevich,
%\emph{Intersection theory on the moduli space of curves},
Funk. Anal. Prilozh., \textbf{25:2} (1991) 50-57;
%\emph{Intersection theory on the moduli space of curves
%and the Airy function},
Comm.Math.Phys. {\bf 147} (1992) 1-23; \\
M.Adler and P. van Moerbeke, Comm.Math.Phys. {\bf 147} (1992) 25;\\
S.Kharchev, A.Marshakov, A.Mironov, A.Morozov and A.Zabrodin,
%\emph{Unification of All String Models with $c<1$},
Phys. Lett. \textbf{B275} (1992) 311-314, hep-th/9111037;
%\emph{Towards unified theory of $2d$ gravity},
Nucl.Phys. \textbf{B380} (1992) 181-240, hep-th/9201013;
%{\it Generalized Kontsevich Model Versus Toda Hierarchy
%and Discrete Matrix Models},
Nucl.Phys. \textbf{B397} (1993) 339-378, hep-th/9203043;
%\emph{Landau-Ginzburg Topological Theories in the Framework of GKM and Equivalent Hierarchies},
Mod.Phys.Lett. \textbf{A8} (1993) 1047-1062,
Theor. Math. Phys. \textbf{95} (1993) 571-582, hep-th/9208046; \\
A.Marshakov, A.Mironov and A.Morozov,
%\emph{From Virasoro Constraints in Kontsevich's Model to
%$\cal W$-constraints in 2-matrix Models},
Mod. Phys. Lett. A7 (1992) 1345-1360, hep-th/9201010;
%\emph{On the Equivalence of Topological and
%Quantum 2d Gravity},
Phys. Lett. \textbf{274B} (1992) 280-288, hep-th/9201011;\\
P.Di Francesco, C.Itzykson and J.-B.Zuber,
%\emph{Polynomial averages in the Kontsevich model},
Comm.Math.Phys. {\bf 151} (1993) 193-219, hep-th/9206090

\bibitem{Kle}
A.Klemm and P.Sulkowski,
\emph{Seiberg-Witten theory and matrix models},
arXiv:0810.4944; \\
P.Sulkowski,
\emph{Matrix models for 2* theories},
arXiv:0904.3064

\bibitem{Nek}
N.Nekrasov and A.Okounkov,
%{\it Seiberg-Witten Theory and Random Partitions},
arXiv:hep-th/0306238; \\
A.Marshakov and N.Nekrasov,
%\emph{Extended Seiberg-Witten Theory and Integrable Hierarchy},
JHEP \textbf{0701} (2007) 104, hep-th/0612019

\bibitem{EyKle}
B.Eynard,
\emph{All order asymptotic expansion of large partitions},
arXiv:0804.0381

\bibitem{unint}
A.Morozov,
\emph{Unitary Integrals and Related Matrix Models},
arXiv:0906.3518

\bibitem{EyFT}
B.Eynard,
%\emph{All genus correlation functions for the hermitian 1-matrix model},
JHEP \textbf{0411} (2004) 031, hep-th/0407261; \\
B.Eynard and N.Orantin,
%\emph{Topological expansion of the 2-matrix model correlation functions:
%diagrammatic rules for a residue formula},
JHEP \textbf{0612} (2006) 026, math-ph/0504058;
%{\it Invariants of algebraic curves and topological expansion},
math-phys/0702045; \\
L.Chekhov, B.Eynard,
%\emph{Hermitian matrix model free energy:
%Feynman graph technique for all genera},
JHEP \textbf{0603} (2006) 014, hep-th/0504116;
%\emph{Matrix eigenvalue model: Feynman graph technique for all genera},
JHEP \textbf{0612} (2006) 026, math-ph/0604014; \\
B.Eynard, M.Marino and N.Orantin,
%\emph{Holomorphic anomaly and matrix models},
JHEP \textbf{0706} (2007) 058, hep-th/0702110; \\
N.Orantin, PhD thesis, arXiv:0709.2992;
%\emph{Gaussian matrix model
%in an external field and non-intersecting Brownian motions},
arXiv:0803.0705


\bibitem{AMM}
A.Alexandrov, A.Mironov and A.Morozov,
%{\it Partition functions of matrix models as the first special
%functions of string theory. I. Finite-size Hermitian
%1-matrix model},
Int.J.Mod.Phys. {\bf A19} (2004) 4127,
Theor.Math.Phys. {\bf 142} (2005) 349, hep-th/0310113;
%{\it Unified description of correlators in non-Gaussian
%phases of Hermitian matrix model},
Int.J.Mod.Phys.
{\bf A21} (2006) 2481, hep-th/0412099;
%{\it Solving Virasoro constraints in matrix models},
Fortsch. Phys. {\bf 53} (2005) 512, hep-th/0412205;\\
A.Alexandrov, A.Mironov, A.Morozov and P.Putrov,
%{\it Partition functions of matrix models as the first special
%functions of string theory. II. Kontsevich model},
Int.J.Mod.Phys. {\bf A} (2009), arXiv:0811.2825

\end{thebibliography}

\documentstyle[amsmath,amssymb,graphicx]{article}

\def\be{\begin{eqnarray}}
\def\ee{\end{eqnarray}}
\def\nn{\nonumber}

\def\tr{{\rm tr}\,}

\hoffset= - 1.0in         % switch off for draft style
\begin{document}

\hfill ITEP/TH-24/09

\bigskip

\centerline{\Large{On Equivalence of two Hurwitz Matrix Models
}}

\bigskip
\begin{center}
%\vspace{0.3cm}
%\emph{} }\\
{\large A.Morozov and Sh.Shakirov}\\
\vspace{0.4cm}
{\em ITEP, Moscow, Russia\\
MIPT, Dolgoprudny, Russia}\\
\end{center}

\bigskip

\centerline{ABSTRACT}

\bigskip

{\footnotesize
In
%arXiv:0902.2627
\cite{Wops}
a matrix model representation was found
for the simplest Hurwitz partition function,
which has Lambert curve $\phi e^{-\phi} = \psi$
as a classical equation of motion.
We demonstrate that Fourier-Laplace transform
in the logarithm of external field $\Psi$
converts it into a more sophisticated form,
recently suggested in
%arXiv:0906.1206.
\cite{BEMS}.
}

\bigskip

\bigskip

Representation of Riemann surfaces (curves) as ramified coverings
of some simpler (bare) surfaces, like Riemann sphere,
is a very important technique in string theory,
used from summation of perturbation theory series \textit{a la}
\cite{KniUFN,UFN2} to description of partition functions
as $\tau$-functions of integrable hierarchies \cite{intau}-\cite{DV}.
For many years, however, this subject looked too difficult to
be addressed in a systematic way, and only recently the
underlying theory of Hurwitz-Kontsevich (HK) partition functions
-- concise language for accumulating information about
ramified coverings -- began to attract attention.
Fortunately, now activity in this field is fastly developing,
see \cite{Hur} for some important references.
This theory is only in its very beginning and it is, perhaps,
too early to apply it to the above-mentioned old problems
of general value
(of which the main one is development of the free-field calculus
{\it a la} \cite{KniUFN,fref} on arbitrary coverings),
but it is quite interesting and puzzling by itself.
Being related to explicitly topological theories,
the HK partition functions should of course possess matrix-model
representations, and their place in the "M-theory of matrix models"
\cite{Mth} is one of the first problems to be addressed in this area.

Recently, in the simplest example of "ordinary" Hurwitz numbers
such matrix model was explicitly derived in \cite{Wops}.
It belongs to the family of Generalized Kontsevich models
\cite{GKM,UFN3}, but has a somewhat unusual form:

\begin{equation}
\addtolength{\fboxsep}{5pt}
\boxed{
\begin{gathered}
Z_{HK}(\Psi) \sim \int_{n\times n}
d \mu(\phi) \
\exp\left( - \dfrac{1}{2g} \tr \phi^2 +
\tr \big( e^{\phi - ng/2} \Psi \big) \right) \end{gathered}
}\label{HarerZagierDuo}
\end{equation}
\smallskip\\
where $\Psi$ is an $n\times n$ matrix-valued Miwa variable, through which
the ordinary time-variables are expressed as
$p_k = kt_k = \tr \Psi^k$, and the integral is taken over all Hermitian $n \times n$ matrices $\phi$ with the measure

\begin{align}
d\mu(\phi) = \sqrt{{\det} \left(
\dfrac{ \sinh \left(\frac{\phi\otimes I - I\otimes\phi}{2}\right)}
{ \left(\frac{\phi\otimes I - I\otimes\phi}{2}\right)} \right)} d\phi
\ =\  \Delta(\phi)\Delta(e^{\phi})\
[dU] \prod\limits_{i=1}^n e^{-\frac{n-1}{2} \phi_i} d\phi_i
\label{meas}
\end{align}
\smallskip\\
This measure is one of the main peculiarities of the model.
It is in fact nicely matched with the exponential coupling to
external field $\Psi$ in the action, so that trigonometric
Van-der-Monde determinant
$\Delta(e^\phi) = \prod_{i<j} (e^{\phi_i}-e^{\phi_j})$
is absorbed into the Itzykson-Zuber integral over unitary matrices $U$.
Another attractive feature is that in neglect of this
measure the classical equation of motion has the desired
form of Lambert curve,
\begin{align}
\phi e^{-\phi} = e^{-ng/2}\psi,
\end{align}
which is well known from many papers of ref.\cite{Hur}
to play an important role in the theory of Hurwitz numbers.
Last week an alternative matrix model
representation was suggested in \cite{BEMS}
for the {\it same} partition function,
which looks more sophisticated and is very much in the spirit of \cite{Kle}
(that paper is in turn a development of \cite{Nek} and especially of \cite{EyKle}).
This representation is rather natural from the perspective of character
representations from \cite{Wops} and \cite{unint}, and is also closely related
to the matrix-model field theory of \cite{EyFT}
(a parallel development to \cite{AMM,Mth}).

The goal of this short note is to describe the explicit
relation between these two currently-available matrix model
representations of the simplest HK partition function:
one is nothing but the Fourier-Laplace transform of another.
This means that the pertinent advantages and relations of
each of these two models are immediately inherited by the
other, what can open additional possibilities for further
developments. The most important could be generalizations
to multiple and open Hurwitz numbers and a progress in
understanding of the puzzling Virasoro constraints and
related link between the HK partition function and conventional
cubic Kontsevich theory -- the main building block of the
M-theory of matrix models. See the last few papers in \cite{Hur}
for clearer formulation of these problems.

\bigskip

We want to find the shape of potential $V(M)$ from

\be
\boxed{
Z_{HK}\Big(\Psi = e^R\Big) = Z_{BEMS}(R),
}
\label{claim}
\ee
where

\begin{align}
Z_{HK}(\Psi) \ = \ & g^{-n^2/2} \exp\left(- g \dfrac{n(n^2-1)}{24} \right) \int
d \mu(\phi) \
\exp\left( - \dfrac{1}{2g} \tr \phi^2 +
\tr \big( e^{\phi - ng/2} \Psi \big) \right) = \nn \\ &
\nonumber \\ &
= g^{-n^2/2} \exp\left( - g \dfrac{n^3}{6} + g \dfrac{n}{24} \right) \int \dfrac{\Delta(\phi)}{\Delta(\Psi)} \prod\limits_{i = 1}^{n}
\exp\left( - \dfrac{\phi^2_i}{2g} - \dfrac{2n-1}{2} \phi_i + e^{\phi_i} \psi_i \right) d\phi_i
\label{zhk}
\end{align}
and

\begin{align}
Z_{BEMS}(R) \ = \ & g^{-n^2} \dfrac{\Delta( R )}{\Delta(e^{R})} \int dM \exp\left(-\dfrac{1}{g} \tr V(M) + \dfrac{1}{g} \tr MR \right)
= \label{zbemsx} \\ & \nonumber \\ & = g^{-n^2} \int \dfrac{\Delta(M)}{\Delta( e^{R} )}
\prod\limits_{i = 1}^{n} \exp\left( - \dfrac{m_i r_i - V(m_i)}{g} \right) dm_i
\label{zbems}
\end{align}
\smallskip\\
The integrals are taken over Hermitian $n \times n$ matrices $\phi, M$ and $R$ with eigenvalues $\phi_i, m_i$ and $r_i$, respectively. As usual, $\Delta(\phi) = \prod_{i < j} \big(\phi_i - \phi_j\big)$ and
$\Delta(e^\phi) = \prod_{i<j} (e^{\phi_i}-e^{\phi_j})$
are Van-der-Monde determinants. Note that, the $n$-dependent prefactors in the first and second lines of (\ref{zhk}) are different, because in the second line a convenient shift $\phi \mapsto \phi + ng/2$ is made and because of the exponential prefactor from (\ref{meas}), see \cite{Wops} for more details. Note also, that the $R$-dependent pre-factor in (\ref{zbemsx}) is important for
potential $V(M)$ to exist. According to (\ref{claim})-(\ref{zbems}) this
$V(M)$ is given by inverse Fourier-Laplace transform,

\be
e^{- \tr V(M)/g } = \int dR \dfrac{\Delta(e^R)}{\Delta(R)} \exp\left( -\dfrac{1}{g} \tr MR \right) Z_{HK} \big(e^R\big)
\ee
\smallskip\\
Note that $g^{-n^2}$ disappears after the transform. Substituting the expression for $Z_{HK}\big(e^R\big)$, we obtain

\begin{align*}
e^{- \tr V(M)/g } = g^{-n^2/2} \int dR \dfrac{\Delta(e^R)}{\Delta(R)} \exp\left( -\dfrac{1}{g} \tr MR \right) \int
d \mu(\phi) \ \exp\left( - \dfrac{1}{2g} \tr \phi^2 +
\tr \big( e^{\phi - ng/2} e^R \big) - g \dfrac{n(n^2-1)}{24}\right)
\end{align*}
The simplest option now is to diagonalize
$\phi = U \phi_{\rm diag} U^{-1},
R = {\widetilde U} R_{\rm diag} {\widetilde U}^{-1}$
and take Itzykson-Zuber integrals

\be
\int dU \exp (\tr UXU^\dagger Y) = \frac{\det_{ij} e^{x_iy_j}}{\Delta(X)\Delta(Y)}
\ee
\smallskip\\
over $U$ and ${\widetilde U}$. What remains is a pair of eigenvalue integrals:

\begin{align*}
e^{- \tr V(M)/g } = e^{ - g n^3/6 + gn/24} \int \prod\limits_{i = 1}^{n} \exp\left( - \dfrac{1}{g} m_i r_i \right) dr_i
\int \dfrac{g^{-n^2/2} \Delta(\phi)}{\Delta(M)}
\prod\limits_{i = 1}^{n} \exp\left( - \dfrac{\phi^2_i}{2g} - \dfrac{2n-1}{2} \phi_i + e^{\phi_i + r_i}
\right) d \phi_i
\end{align*}
Shifting $r_i \mapsto r_i - \phi_i$, we obtain

\begin{align*}
e^{- \tr V(M)/g } = e^{ - g n^3/6 + gn/24} \int \prod\limits_{i = 1}^{n} \exp\left( e^{r_i} - \dfrac{1}{g} m_i r_i \right) dr_i
\int \dfrac{ g^{-n^2/2} \Delta(\phi)}{\Delta(M)}
\prod\limits_{i = 1}^{n} \exp\left( - \dfrac{\phi^2_i}{2g} + \dfrac{1}{g} C_i \phi_i \right) d \phi_i
\end{align*}
where $C_i = m_i - g \dfrac{2n-1}{2} $. Using the Itzykson-Zuber formula and the Gaussian integral, it is easy to show that

\be
\int \Delta(\phi) \prod\limits_{i = 1}^{n} \exp\left( - \dfrac{\phi^2_i}{2g} + \dfrac{1}{g} C_i \phi_i \right) d \phi_i
= g^{n/2} \Delta(C) \prod\limits_{i = 1}^{n} \exp\left( \dfrac{C_i^2}{2g} \right)
\ee
Consequently,

\be
e^{- \tr V(M)/g } = g^{-n^2/2+n/2} \exp\left( - g\dfrac{n^3}{6} + g\dfrac{n}{24} + \dfrac{1}{2g} \sum\limits_{i = 1}^{n} \Big( m_i - g \dfrac{2n - 1}{2} \Big)^2 \right) \ \prod\limits_{i = 1}^{n} \int dr_i \exp\left( e^{r_i} - \dfrac{1}{g} m_i r_i \right)
\ee
\smallskip\\
The last integral can be taken explicitly by making a change of variable $\psi = e^r$:

\be
 f(m) = \int dr \exp\left( e^r - \dfrac{m}{g} r \right)
 = \int \dfrac{d\psi}{\psi} \exp\left( \psi - \dfrac{m}{g} \log \psi \right)
 = \int d\psi \ \psi^{- m/g - 1} e^{\psi} = e^{-i\pi m/g}\Gamma\big( - \dfrac{m}{g}\big)
\ee
\smallskip\\
The last equality is nothing but the integral representation of
(appropriate analytic continuation of) Euler's $\Gamma$-function.
It is, of course, correctly defined only for certain choice of integration contour.
We finally obtain

\begin{align}
e^{- \tr V(M)/g } = \prod\limits_{j = 1}^{n} e^{-V(m_j)/g} = \prod\limits_{j = 1}^{n} g^{(1-n)/2} \exp\left( - g\dfrac{n^2}{6} + \dfrac{g}{24} + \dfrac{1}{2g}
\Big( m_j - g \dfrac{2n - 1}{2} \Big)^2 - i \pi \dfrac{m_j}{g} \right)\Gamma\left(-\dfrac{m_j}{g}\right)
\end{align}
\smallskip\\
where the r.h.s. is indeed a product over all eigenvalues of $M$. Thus, the potential $V(M)$ exists and is equal to

\be
V(x) = i \pi x - g \log \Gamma\left(-\dfrac{x}{g}\right) + g^2\dfrac{n^2}{6} - \dfrac{g^2}{24} - \dfrac{1}{2}
\Big( x - g\dfrac{2n - 1}{2} \Big)^2 + \dfrac{n-1}{2} g \log g
\ee
\smallskip\\
or, after opening the brackets,

\be
\boxed{
V(x) = - \dfrac{x^2}{2} - g \log \Gamma\left(-\dfrac{x}{g}\right) + i \pi x + \left( n - \dfrac{1}{2} \right) g x + \dfrac{n-1}{2} g \log g  - \dfrac{g^2}{3} \left( n^2 - \dfrac{3}{2}n + \dfrac{1}{2} \right)
}
\label{result}
\ee
\smallskip\\
This expression coincides with eq.(8) of \cite{BEMS}. Note only that the very last term of this formula
in \cite{BEMS} contains $2$ instead of $1/2$, because of a misprint
in original expression for $f_\lambda(C_2)$ in that paper.
The right formula for $f_\lambda(C_2)$ is \linebreak
\be
f_\lambda(C_2) =
\frac{1}{2}\sum_{i=1}^n \lambda_i(\lambda_i-2i+1) =
\frac{1}{2}\sum_i h_i^2 - \left(n-\frac{1}{2}\right)\sum_i h_i
+ \frac{n}{3}\left(n^2-\frac{3}{2}n+\frac{1}{2}\right),
\ee
where $h_i = \lambda_i - i + n$, see eq.(26) in \cite{Wops}.
There is one more term in $V(x)$ in \cite{BEMS}, which is not present in
(\ref{result}): namely, $x \log g$.
It results from a slightly different definition of the HK partition function
in \cite{BEMS}:
it differs from that of \cite{Wops} by rescaling of the external field
$\Psi \rightarrow \Psi/g$.
This shifts $\ \tr MR = \tr M\log\Psi\ $ by $\ \log(1/g) \tr M$, what
exactly implies adding $x \log g$ to $V(x)$.

We conclude, that the two matrix models --
the model of \cite{BEMS} with the logarithm-of-a-gamma-function potential
and the model of \cite{Wops} with the trigonometric-Van-der-Monde determinant
-- are in fact equivalent.
Since this equivalence is now made explicit, both models can be used on equal footing.

\newpage

\section*{Acknowledgements}

Our work is partly supported by Russian Federal Nuclear Energy
Agency, by the joint grants 09-02-91005-ANF, 09-02-90493-Ukr,
09-01-92440-CE and 09-02-9305-CNRSL,
by the Russian President's Grants of
Support for the Scientific Schools NSh-3035.2008.2
and by the RFBR grant 07-02-00645.

\end{document}